A boat floating in fresh water displaces water weighing 35.0 kN.

(a) What is the weight of the water that this boat displaces when floating in salt water of density 1.10 ✕ 103 kg/m3?
___ kN

(b) What is the difference between the volume of fresh water displaced and the volume of salt water displaced (VfreshVsalt)? (Take the density of fresh water to be 998 kg/m3.)
___m3
Answer #1

Solution of A:

We have,
Weight of the water displaced by floating boat \rightarrow W_{w}=35\ kN=35\times10^{3}\ N
Density of the salt water \rightarrow \rho_{s}=1.10\times10^{3}\ kg\ m^{-3}
Density of the water -3 998 kg m +Pw
Gravitational acceleration +9= 9.81 m/s2

The mass m of the water is calculated as:

m=\frac{W_{w}}{g}

\therefore m=\frac{35\times10^{3}}{9.81}

\therefore m=3567.79\ kg

The volume V of the water is calculated as:

V=\frac{m}{\rho_{s}}

\therefore V=\frac{3567.79}{1.10\times10^{3}}

\therefore V=3.2434\ m^{3}

The weight W_{s} of the salt water is calculated as:

W_{s}=\rho_{s}Vg

\therefore W_{s}=1.40\times10^{3}\times3.2434\times9.81

\therefore W_{s}=44544.8556\ N

{\color{Blue} \therefore W_{s}=44.545\ kN}

Therefore, the displaced weight of the salt water is  {\color{Blue} W_{s}=44.545\ kN}.

Solution of B:

The volume V_{fresh} of the water in fresh water:

W_{w}=\rho_{w}V_{fresh}g

\therefore V_{fresh}=\frac{W_{w}}{\rho_{w}g}

\therefore V_{fresh}=\frac{35\times10^{3}}{998\times9.81}

{\color{Blue} \therefore V_{fresh}=3.5749\ m^{3}}

Difference between the volume of fresh water displaced and the volume of salt water displaced is calculated as:

\therefore V_{fresh}-V=3.5749-3.2434

{\color{Blue} \therefore V_{fresh}-V=0.3315\ m^{3}}

Difference between the volume of fresh water displaced and the volume of salt water displaced is {\color{Blue} 0.3315\ m^{3}} .